the perimeter of an isoscele triangle is 32 the lengths of the two sids are 9 in. less than 2 times the length of the base find the lengths of three sides
Answer by mananth(16946)

the perimeter of an isoscele triangle is 32 the lengths of the two sids are 9 in. less than 2 times the length of the base find the lengths of three sides
..
let base be x
side = 2x-9
perimeter =32
...
x+2x-9+2x-9=32
5x-18=32
5x=50
/5
x=10 the base
..
sides will be 2x-9 = 20-9=11
11+11+10=32
...
m.ananth@hotmail.ca